# Gauss-Newton Natural Gradient Descent for Shape Learning

### v1

@@ 1 Introduction
  • We implement a memory- and time-efficient Gauss-Newton method tailored to geometric learning tasks, enabling scalability to larger network sizes.
@@ 4.1 Plateau’s Problem
In shape optimization, a key problem is learning a smooth surface that connects to a given interface. We focus on learning minimal surfaces as a special case. The problem of finding a minimal surface enclosed by a given Jordan curve is called Plateau’s Problem. Minimal surfaces locally minimize surface area [23], which is equivalent to having zero mean curvature. We consider the catenoid and Enneper surface—two classical solutions with known analytic forms—as test cases. These examples allow us to assess how well different optimization methods minimize mean curvature while fitting to the prescribed interface. For error measurement, we use the Chamfer Divergence $CD_{1}(P,Q)$ between the model surface $P$ and the true surface $Q$ (see Section C). Figures 1 to 4 show the results of training for 20 minutes, in which Gauss-Newton completed $\sim 8000$ iterations, LBFGS $\sim 9000$ and Adam $\sim 17000$. Our results show that Gauss-Newton achieves lower losses, errors, and faster convergence than Adam and LBFGS.
Figure 1: The absolute mean-curvature $\left\lvert\kappa_{H}(\mathbf{x})\right\rvert$ plotted on the best model surface trained with Adam (left), LBFGS (middle), Gauss-Newton (right) for 20 minutes.
Figure 2: Comparison of Loss (interface + mean curvature) and Chamfer Discrepancy Over Time for the Catenoid.
Figure 3: $\left\lvert\kappa_{H}(\mathbf{x})\right\rvert$ plotted on the best model surface trained with Adam (left) and Gauss-Newton (right) for 20 minutes. LBFGS becomes unstable.
Figure 4: Comparison of Loss (interface + mean curvature) and Chamfer Discrepancy Over Time for the Enneper Surface.
@@ 4.2 Developable Surfaces
Another class of geometrically constrained surfaces is developable surfaces, characterised by having zero Gaussian curvature $\kappa_{G}$ at every point. Developable surfaces can be flattened onto a plane without distortion. This constraint appears naturally in applications such as sheet metal design and architectural geometry. To assess performance, we train our model to represent the mantle of a truncated cone, a canonical example of a smooth developable surface. Similar to the minimal surface experiments, we enforce an interface constraint and minimize a geometric objective - here the Gauss-curvature.
@@
(a) Heatmaps of absolute Gauss curvature.
@@
Figure 5: Comparison of Adam and Gauss-Newton optimization on the truncated cone problem. (a) The absolute Gauss curvature $\left\lvert\kappa_{G}(\mathbf{x})\right\rvert$ is plotted for the best model surface trained with Adam (left) and Gauss-Newton (right) for 20 minutes. LBFGS becomes unstable and is omitted. (b) Total loss (interface + Gauss-curvature) over time for the truncated cone.
@@ 4.3 Implicit Neural Shapes
We evaluate our method on two standard benchmark shapes—the Rocker Arm and the Stanford Bunny—using a loss that enforces closeness to the ground truth surface points, encourages alignment with the prescribed normals, and is regularized by an eikonal penalty. As shown in Figures 6(b) and 7(b), Gauss-Newton consistently achieves the lowest loss across both examples, outperforming Adam and LBFGS. The qualitative results in Figures 6(a) and 7(a) further demonstrate that Gauss-Newton produces reconstructions with finer geometric detail and better visual agreement with the target meshes. We also found Gauss-Newton to be particularly effective in data-sparse settings, achieving similar mesh quality and loss reduction when trained on only 20-30% of the available ground truth data. In these experiments, we randomly discarded a large portion of the surface data, using only the remaining points for training. Despite this significant reduction in data, the resulting meshes and loss curves closely matched those from the full data experiments, with the added benefit of faster Gauss-Newton steps. This highlights the usefulness of Gauss-Newton for applications where only limited data is available for reconstruction.
(a) The best surface after 10 minutes of training with Adam, LBFGS, and Gauss-Newton, respectively.
@@
Figure 6: Comparison of different optimizers on the Rocker Arm problem minimizing (interface + normal). Gauss-Newton $\sim 800$ iterations, LBFGS $\sim 6100$, Adam $\sim 34000$.
(a) The best surface after 10 minutes of training with
Adam, LBFGS, and Gauss-Newton, respectively.
@@
Figure 7: Comparison of different optimizers on the Stanford Bunny problem minimizing (interface + normal). Gauss-Newton $\sim 350$ iterations, LBFGS $\sim 800$, Adam $\sim 28000$.
@@ 4.4 Jet Engine Bracket
The last use-case considers the design of a jet engine bracket [15, 4]. Conceptually, this geometry-informed neural network task is similar to the first two experiments, but with more numerous and computationally costly constraints. These include an eikonal regularization, a non-trivial design region enclosed by a watertight mesh, and six cylindrical interfaces with inward-facing normals. Additionally, we minimize surface strain, which is another curvature-based quantity, but different to the previous use-cases, it is not zero everywhere for the (unknown and possibly non-unique) optimal shape. We expect this loss contribution to be severely ill-conditioned.
@@
(a) Heatmaps of surface strain.
@@
Figure 8: Comparison of Gauss-Newton and Adam optimization on the jet engine bracket problem without clipping the ill-conditioned strain. (a) The surface strain $\kappa_{1}^{2}(\mathbf{x})+\kappa_{2}^{2}(\mathbf{x})$ in $\log_{10}$-scale for the shapes identified by Gauss-Newton (left) and Adam (right). Adam fails to converge to a feasible shape. (b) Total loss over time corresponding to the jet engine brackets in (a).
(a) Heatmaps of surface strain.
@@
Figure 9: Comparison of Gauss-Newton and Adam optimization on the jet engine bracket problem with clipping the ill-conditioned strain. (a) The surface strain $\kappa_{1}^{2}(\mathbf{x})+\kappa_{2}^{2}(\mathbf{x})$ in $\log_{10}$-scale for the shapes identified by Gauss-Newton (left) and Adam (right). (b) Total loss over time. Additionally plotting the ill-conditioned strain contribution reveals that Adam fails to optimize it at all. While clipping helps Adam significantly, it is still outperformed by Gauss-Newton.
@@ Appendix C Chamfer Divergence
Appendix D Hyperparameters and Model Architectures

### v2

@@ 1 Introduction
  • We implement111code available at https://github.com/JamesAndrewKing/GaussNewtonShapes a memory- and time-efficient Gauss-Newton method tailored to geometric learning tasks, enabling scalability to larger network sizes.
@@ 4.1 Plateau’s Problem
In shape optimization, a key problem is learning a smooth surface that connects to a given interface. We focus on learning minimal surfaces as a special case. The problem of finding a minimal surface enclosed by a given Jordan curve is called Plateau’s Problem. Minimal surfaces locally minimize surface area [23], which is equivalent to having zero mean curvature. We consider the catenoid and Enneper surface—two classical solutions with known analytic forms—as test cases. These examples allow us to assess how well different optimization methods minimize mean curvature while fitting to the prescribed interface. For error measurement, we use the Chamfer Divergence $CD_{1}(P,Q)$ between the model surface $P$ and the true surface $Q$ (see Section C). Figures 1(a) to 2(b) show the results of training for 20 minutes, in which Gauss-Newton completed $\sim 8000$ iterations, LBFGS $\sim 9000$ and Adam $\sim 17000$. Our results show that Gauss-Newton achieves lower losses, errors, and faster convergence than Adam and LBFGS.
Refer to captionAdamLBFGSGauss-Newton
(a) Absolute mean curvature $\left\lvert\kappa_{H}(\mathbf{x})\right\rvert$ of the best surface.
(b) Loss over time.
Figure 1: Plateau’s problem on the catenoid geometry with mean curvature and interface loss. The prescribed interface is highlighted in green in (a). The chamfer divergence to the ground truth is additionally shown in Figure 8(a).
Refer to captionAdamGauss-Newton
(a) Absolute mean curvature $\left\lvert\kappa_{H}(\mathbf{x})\right\rvert$ of the best surface.
(b) Loss over time.
Figure 2: Plateau’s problem on the Enneper minimal surface geometry with mean curvature and interface loss. The prescribed interface is highlighted in green in (a). LBFGS becomes unstable and is omitted. The chamfer divergence to the ground truth is additionally shown in Figure 8(b).
@@ 4.2 Developable Surfaces
Another class of geometrically constrained surfaces is developable surfaces, characterised by having zero Gaussian curvature $\kappa_{G}$ at every point. Developable surfaces can be flattened onto a plane without distortion. This constraint appears naturally in applications such as sheet metal design and architectural geometry. To assess performance, we train our model to represent the mantle of a truncated cone, a canonical example of a smooth developable surface. Similar to the minimal surface experiments, we enforce an interface constraint and minimize a geometric objective - here the Gauss curvature.
@@
Refer to captionAdamGauss-Newton
(a) Absolute Gauss curvature $\left\lvert\kappa_{G}(\mathbf{x})\right\rvert$ of the best surface.
@@
Figure 3: Developable surface problem on the truncated cone geometry with Gauss curvature and interface loss. The prescribed interface is highlighted in green in (a). LBFGS becomes unstable and is omitted.
@@ 4.3 Implicit Neural Shapes
We evaluate our method on two standard benchmark shapes—the Rocker Arm and the Stanford Bunny—using a loss that enforces closeness to the ground truth surface points, encourages alignment with the prescribed normals, and is regularized by an eikonal penalty. As shown in Figures 4(b) and 5(b), Gauss-Newton consistently achieves the lowest loss across both examples, outperforming Adam and LBFGS. The qualitative results in Figures 4(a) and 5(a) further demonstrate that Gauss-Newton produces reconstructions with finer geometric detail and better visual agreement with the target meshes. We also found Gauss-Newton to be particularly effective in data-sparse settings, achieving similar mesh quality and loss reduction when trained on only 20-30% of the available ground truth data. In these experiments, we randomly discarded a large portion of the surface data, using only the remaining points for training. Despite this significant reduction in data, the resulting meshes and loss curves closely matched those from the full data experiments, with the added benefit of faster Gauss-Newton steps. This highlights the usefulness of Gauss-Newton for applications where only limited data is available for reconstruction.
Refer to captionAdamLBFGSGauss-Newton
(a) The best surface.
@@
Figure 4: Implicit neural shape problem on the rocker arm geometry with interface and normal loss.
Refer to captionAdamLBFGSGauss-Newton
(a) The best surface.
@@
Figure 5: Implicit neural shape problem on the Stanford bunny geometry with interface and normal loss.
@@ 4.4 Jet Engine Bracket
The last use-case considers the design of a jet engine bracket [15, 4]. Conceptually, this geometry-informed neural network task is similar to the first two experiments, but with more numerous and computationally costly constraints. These include an eikonal regularization, a non-trivial design region enclosed by a watertight mesh, and six cylindrical interfaces with inward-facing normals illustrated in Figure 9. Additionally, we minimize surface strain, which is another curvature-based quantity, but different to the previous use-cases, it is not zero everywhere for the (unknown and possibly non-unique) optimal shape. We expect this loss contribution to be severely ill-conditioned.
@@
Refer to captionGauss-NewtonAdam
(a) Surface strain of the best surface.
@@
Figure 6: Jet engine bracket problem without clipping the ill-conditioned strain. The additional constraints include interface, interface normal, design region, connectedness, and eikonal regularization, illustrated in Figure 9. (a) The surface strain $\kappa_{1}^{2}(\mathbf{x})+\kappa_{2}^{2}(\mathbf{x})$ in $\log_{10}$-scale for the best shapes identified by each optimizer. Adam fails to converge to a feasible shape as also quantified by the total loss over time in (b).
Refer to captionGauss-NewtonAdam
(a) Surface strain of the best surface.
@@
Figure 7: Jet engine bracket problem with clipping the ill-conditioned strain. The additional constraints include interface, interface normal, design region, connectedness, and eikonal regularization, illustrated in Figure 9. (a) The surface strain $\kappa_{1}^{2}(\mathbf{x})+\kappa_{2}^{2}(\mathbf{x})$ in $\log_{10}$-scale for the best shapes identified by each optimizer. (b) Total loss over time. Additionally plotting the ill-conditioned strain contribution reveals that Adam fails to optimize the strain at all. While clipping stabilizes and helps Adam significantly, it is still outperformed by Gauss-Newton.
@@ Appendix C Chamfer Divergence
(a) Catenoid (cf. Figure 1).
(b) Enneper surface (cf. Figure 2).
Figure 8: Chamfer divergences over time quantifying the deviation from the known ground truth. These relate similarly to the training loss curves with Gauss-Newton converging both faster and to lower values.
Appendix D Jet Engine Bracket Constraints Visualized
Figure 9: The constraints of the jet engine bracket problem illustrated. Any generated shape must be contained in the sketched design region and attach to the six cylindrical interfaces in green.
Appendix E Hyperparameters and Model Architectures